OpenStudy (kainui):

Conjecture: 5 is the only prime that is the sum of all preceding primes.

ganeshie8 (ganeshie8):

Interesting... The difference between two consecutive prime numbers appears to mostly increase for large numbers. So there might exist a large prime that is sum of all the preceding primes.

ganeshie8 (ganeshie8):

2 + 3 = 5 2 + 3 + 5 = 10 ** 2 + 3 + 5 + 7 = 17 2 + 3 + 5 + 7 + 11 = 28 ** \(p_{2n}\) cannot be the sum of its preceding primes as the sum is even. So 50% of the primes are in favor of your conjecture ! :)

OpenStudy (kainui):

Haha nice good thinking!

ganeshie8 (ganeshie8):

How strongly do you believe your conjecture is true ? (I'm looking at the lower/upper bounds for the primes - might give us something to mess with...)

ganeshie8 (ganeshie8):

Nvm, twin primes kill that idea of mine

OpenStudy (kainui):

Yeah this is sorta related to prime gaps so I don't know if it's different enough or not to be provable. I think it might but not sure, I think perhaps there existing at least one prime between powers of 2 could help?

ganeshie8 (ganeshie8):

Definitely, bertrand's postulate right ? https://en.wikipedia.org/wiki/Bertrand%27s_postulate

ganeshie8 (ganeshie8):

\(p_{n+1} \le 2^{2^{n}}\)

ganeshie8 (ganeshie8):

\[S_n = \sum\limits_{k=1}^n p_k \le \sum\limits_{k=1}^n 2^{2^{k-1}}\]

ganeshie8 (ganeshie8):

I think we're done if we could show that sum is less than \(p_{n+1}\)

OpenStudy (kainui):

Hmm that seems difficult I'm not sure if this will end up working out.

ganeshie8 (ganeshie8):

Yeah, I don't really know. I'm just typing whatever comes in my head haha

OpenStudy (kainui):

I'm feeling that I thought this was gonna work out too but it's sorta like bounded on the wrong side of the fence. I am looking at this now: \[p_{n+1} < \sum_{k=1}^n p_k\] I think this slightly stronger statement is true for \(n>3\).

ganeshie8 (ganeshie8):

Ahh right, the sum might exceed our fancy prime; so we should be looking at the lower bound

imqwerty (imqwerty):

Maybe we can use that formula for nth prime which ganeshie once posted.. That won't be easy tho

OpenStudy (kainui):

I think I might have almost proved it! Bertrand's postulate: \[p_{k+1} < 2 p_k\]\[p_{k+1}-p_k < p_k\]\[\Delta p_k < p_k\]\[\sum_{k=1}^n \Delta p_k < \sum_{k=1}^n p_k\]\[p_{n+1}-2 < \sum_{k=1}^n p_k\] which is almost \[p_{n+1} < \sum_{k=1}^n p_k \]

OpenStudy (kainui):

Ah ok I have found the last piece of the puzzle, it's actually really amazing how perfectly it fits, @wio gave me the last info I needed! Instead of summing from 1, I start at some variable. \[\sum_{k=a}^n \Delta p_k < \sum_{k=a}^n p_k\]\[p_{n+1}-p_a < \sum_{k=a}^n p_k\]\[p_{n+1}< p_a+ \sum_{k=a}^n p_k\] Now since I already know \(5=3+2\) or in other words: \(p_3=p_2+p_1\) I plug in \(a=3\) and get: \[p_{n+1}< \sum_{k=1}^2p_k+ \sum_{k=3}^n p_k\] \[p_{n+1}< \sum_{k=1}^n p_k\] Uncannily exactly what I was seeking to prove... Spooky how well that worked out!

ganeshie8 (ganeshie8):

Very clever use of the telescoping sum !